If you deposit $5,000 in an account that pays 5% interest compounded annually, how much money will you have in your account at the end of 15 years? Write an exponential function that represents this situation.

Answer by rfer(16322)

A=P(1+r)^15
A=5000(1.05)^15
A=5000*2.07893
A=$10,394.64